OpenStudy (anonymous):

What is the solution to the equation -2n-2/3=12 n=?

OpenStudy (anonymous):

So you have \[-2n-\frac{ 2 }{ 3 }=12\]

OpenStudy (anonymous):

What's your first step, if you want to isolate n

OpenStudy (anonymous):

-2n-2 over 3 all of its over 3

OpenStudy (anonymous):

Crossing out -2 2 and 3

OpenStudy (anonymous):

Plz help

OpenStudy (anonymous):

Oh ok, it's all over three, so in that case you have\[\frac{ -2n-2 }{ 3 }=12\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So you have to get rid of the three on the bottom before you can add or subtract anything, so how can you get rid of the three?

OpenStudy (anonymous):

Put it on the other side

OpenStudy (anonymous):

but its not addition or subtraction so you have to do something to the whole equation to get rid of that 3

OpenStudy (anonymous):

And how do you do that?

OpenStudy (anonymous):

Crossbit out and divide

OpenStudy (anonymous):

Cross it

OpenStudy (anonymous):

Well it's dividing right now, so how could you decide?

OpenStudy (anonymous):

multiply both sides by 3

OpenStudy (anonymous):

Umm @zakiullah we don't give out answers like that, read the code of conduct

OpenStudy (anonymous):

Crossbit out and divide

jhonyy9 (jhonyy9):

-2n-2 12 ------ = ---- so cross multiplie and will get 3 1 -2n-2 = 3*12 -2n-2 = 36 -2n = 38 divide both sides by (-2) n = 38/(-2) = - 19 n= - 19